OpenStudy (anonymous):

Solve the equation: 16x^2+25y^2=544 (i) xy=12 (ii)

OpenStudy (anonymous):

\[(4x + 5y)^2 = 16x^{2} + 25y^{2} + 2(4x)(5y) = 544 + 40*12 = 1024\] \[ 4x + 5y = \pm32\]

OpenStudy (anonymous):

I x = 5n+2, y = -4(n+2), n element Z x = 5 n+3, y = -4 (n-1), n element Z

OpenStudy (anonymous):

y = -4/5 (x-8), y = -4/5 (x+8)